# Overcoming the Thermal-Noise Limit of Room-Temperature Microwave Measurements by “Cavity Pre-cooling” with a Low-Noise Amplifier —Application to Time-resolved Electron Paramagnetic Resonance

### v1

@@ A.2.1 Number of photons inside a microwave cavity
Figure 9: The cavity pre-cooling technique involves electromagnetically coupling a resonator, that is physically heat-sunk to a room-temperature environment (here referred to as the “hot bath”), to an active cold noise source or “load”, here functioning as a “cold bath”.
A.2.2 Mode temperature inside the cavity
@@
Upon achieving dynamic equilibrium, $\dot{q}=0$, the rate equation in Eq. 3 can thereupon be used to calculate the expected photon number. Defining the mode’s equivalent temperature as $T_{\text{mode}}=\frac{q}{\epsilon}=\frac{qhf_{\text{mode}}}{k_{\text{B}}}$, we obtain, for a single coupling port:
$T_{\text{mode}}=\frac{T_{0}+\kappa T_{\text{cold}}}{1+\kappa}.$ (5)
@@
In words: the mode’s expected temperature is the weighted average temperature of the two baths with which the mode interacts, the weighting of each bath being proportional to its interaction strength with the mode, i.e. its “coupling”. Here, the coupling of the mode to the cavity’s own intrinsic heat bath is (by definition/convention) unity. Fig. 10 displays the mode temperature predicted by Eq. 5. When the coupling factor of the over-coupling loop $\kappa$ is large, the mode temperature $T_{\text{mode}}$ will approach that of the cold load $T_{\text{cold}}$.
@@
Figure 10: Simulated equivalent temperature of an electromagnetic mode inside the cavity as a function of the coupling factor $\kappa$ (horizontal coordinate) and the noise temperature of the cold load $T_{\text{cold}}$ coupled to this mode (vertical coordinate).
A.2.3 Cavity cooling via a lossy transmission line
@@
The input of a low-noise amplifier (LNA) can serve as an active cold load for reducing the thermal noise inside a cavity that is coupled to it. Here we analyse the additional thermal noise generated by a lossy transmission line, such as a length of coaxial cable, or any other lossy component inserted between the cold load and the cavity, as shown generically in Fig. 11. See Chapter 8 of Ref. [36] for the conceptual foundations and mathematics of the “wave approach”[52] used here.
@@
Figure 11: Calculating how the cooling LNA’s noise temperature is degraded by transport along a length of (warm) lossy transmission line.
@@
$P_{l}=LP_{\text{cold}}+(1-L)P_{\text{line}},$ (6)
@@
$P_{l}=(1-\frac{L_{\text{dB}}}{4.34})P_{\text{cold}}+\frac{L_{\text{dB}}}{4.34}P_{\text{line}}.$ (7)
Note that this (slightly more complicated) approximation is more accurate than equation (8-2-5) in Siegman’s book [36]. Based on the relation between noise power and noise temperature, namely $P=k_{\text{B}}TB$, the effective noise temperature of the cold load, upon including the additional noise from a lossy transmission line connected to it, can thereupon be approximated as:
$T_{l}=(1-\frac{L_{\text{dB}}}{4.34})T_{\text{cold}}+\frac{L_{\text{dB}}}{4.34}T_{0},$ (8)
where $T_{\text{cold}}$ the temperature of the cold load itself and $T_{0}=T_{\text{line}}$ is the (here assumed to be room) temperature of the transmission line. Combining Eq. 5 and Eq. 8, the mode’s noise temperature inside the microwave cavity is now given by:
$T^{\prime}_{\text{mode}}=\frac{T_{0}+\kappa[(1-\frac{L_{\text{dB}}}{4.34})T_{\text{cold}}+\frac{L_{\text{dB}}}{4.34}T_{0}]}{1+\kappa}=T_{\text{mode}}^{\text{orig.}}+\Delta T_{\text{link}},$ (9)
where $T_{\text{mode}}^{\text{orig.}}$ is the original mode temperature as given by Eq. 5 above and $\Delta T_{\text{link}}=\frac{L_{\text{dB}}}{4.34}\kappa(T_{0}-T_{\text{cold}})/(1+\kappa)$ is the added noise temperature caused by the lossy transmission line.
A.2.4 Multiple cooling ports to a microwave cavity
We now consider a microwave cavity with two ports, one for cooling, the other for read-out, connected to separate LNAs functioning as cold loads. With the setup depicted in Fig. 12 and applying Eq. 4, the rate equation governing the microwave cavity’s photon number is:
@@ A.2.4 Multiple cooling ports to a microwave cavity
where $\kappa_{over}$ and $\kappa_{crit}$ ($\kappa_{crit}$ = 1) are the coupling factors of the over-coupled and critically-coupled loops, respectively. Analogous to the simpler derivation in the previous mode-temperature analysis above, in dynamic equilibrium when $\dot{q}=0$, the depth of the obtainable cavity cooling can be deduced. The mode temperature $T_{\text{mode}}$ of the system, shown in Fig. 12, is given by:
$T_{\text{mode}}=\frac{T_{0}+(\kappa_{\text{over}}+\kappa_{\text{crit}})T_{\text{cold}}}{1+\kappa_{\text{over}}+\kappa_{\text{crit}}}=\frac{T_{0}+(1+\kappa_{\text{over}})T_{\text{cold}}}{2+\kappa_{\text{over}}}.$ (11)
@@
$T_{\text{mode}}=\frac{T_{0}+\sum_{i=1}^{n}\kappa_{i}T_{i}}{1+\sum_{i=1}^{n}\kappa_{i}},$ (12)
@@
Figure 12: Cavity cooling with two separate cold baths.
A.2.5 Multiple cooling ports to a microwave cavity with lossy transmission lines
In this section, we consider a microwave cavity with multiple cooling ports and lossy transmission lines, integrating concepts from sections 2.3 and 2.4. Combining equations Eq. 8 and Eq. 12, yields
$T_{\text{mode}}=\frac{T_{0}+\kappa_{1}[(1-\frac{L_{\text{dB,1}}}{4.34})T_{\text{cold}}+\frac{L_{\text{dB,1}}}{4.34}T_{0}]+\kappa_{2}[(1-\frac{L_{\text{dB,2}}}{4.34})T_{\text{cold}}+\frac{L_{\text{dB,2}}}{4.34}T_{0}]}{1+\kappa_{1}+\kappa_{2}},$ (13)
where $\kappa_{1}$ and $\kappa_{2}$ represent the coupling coefficients of the first and second cold baths, respectively. We here assume the insertion losses of the transmission lines associated with the cold baths ($L_{\text{dB,1}}$, $L_{\text{dB,2}}$) are small. If however the insertion loss of a connecting cable/device is not small, one needs to revert back to Eq. 6, equivalent to equation in (8-2-4) in Siegman’s book [36]. Based on the relation between noise power and noise temperature, $P=k_{\text{B}}TB$, the equivalent noise temperature of a cold load at temperature $T_{\text{cold}}$, observed through a length of transmission line at temperature $T_{0}$ and of significant linear insertion loss $L=10^{-L_{\text{dB}/10}}$, can be quantified as:
$\begin{split}T_{l}&=LT_{\text{cold}}+(1-L)T_{0}\\ &=10^{-\frac{L_{\text{dB}}}{10}}T_{\text{cold}}+\left(1-10^{-\frac{L_{\text{dB}}}{10}}\right)T_{0}.\end{split}$ (14)
@@ A.2.5 Multiple cooling ports to a microwave cavity with lossy transmission lines
$T_{\text{mode}}=\frac{T_{0}+\kappa_{1}[(1-\frac{L_{\text{dB,1}}}{4.34})T_{\text{cold}}+\frac{L_{\text{dB,1}}}{4.34}T_{0}]+\kappa_{2}[10^{-\frac{L_{\text{dB,2}}}{10}}T_{\text{cold}}+(1-10^{-\frac{L_{\text{dB,2}}}{10}})T_{0}]}{1+\kappa_{1}+\kappa_{2}}.$ (15)
These approximations are analyzed in Fig. 13 below.
@@
Figure 13: How, according to Eq. 14, the forward noise power (in units of absolute temperature $T$) at the far end of a lossy cable depends on the cable’s insertion loss $L_{\text{dB}}$. Here the cable is uniformly maintained at a temperature of $T_{\text{0}}=290$ K throughout its length and the near end of the cable is terminated and cooled by the an ACL whose noise temperature is $T_{\text{cold}}=18.2$ K (dashed cornflower-blue baseline)–corresponding to the input noise temperature of the Qorvo QPL9547EVB-01 LNA used experimentally. The zoomed-in area details how, when the insertion loss is less than 0.5 dB, equation Eq. 8 (red dashed line -“Linear Approx.”), i.e. the linear approximation provided by equation (8-2-5) in Siegman’s book [36], compares against the exact formula Eq. 14 (blue solid line).
A.3 NOISE ANALYSIS
@@
Similar to the previous work by Wu et al. [1], we also adopt the “wave approach” (see Chapter 8 of Siegman [36]) to analyze the flow of noise power from the cavity into our homodyne receiver, as shown in Fig. 14. Equation S8 in the supplemental material of Wu et al.’s paper provides a formula for the factor, $\Delta P$ (in units of dB), by which measured noise power is reduced by mode cooling:
@@
$\Delta P^{\prime}=10\log_{10}\frac{{G_{\text{LNA}}[(T_{\text{min}}+T_{\text{mode}})(1-|\Gamma_{c}|^{2})+4T_{0}\frac{R_{\text{n}}}{Z_{0}}\frac{\lvert\Gamma_{\text{c}}-\Gamma_{\text{opt}}\rvert^{2}}{\lvert 1+\Gamma_{\text{opt}}\rvert^{2}}+T_{\text{image}}]+T_{\text{REC}}}}{{G_{\text{LNA}}[(T_{\text{min}}+T_{\text{mode}}^{0})(1-|\Gamma_{c}^{0}|^{2})+4T_{0}\frac{R_{\text{n}}}{Z_{\text{0}}}\frac{\lvert\Gamma_{c}^{0}-\Gamma_{\text{opt}}\rvert^{2}}{\lvert 1+\Gamma_{\text{opt}}\rvert^{2}}+T_{\text{image}}]+T_{\text{REC}}}},$ (16)
@@
However, this formula needs to be modified slightly for the following reasons: (i) Our current receiver is a homodyne, not a single-conversion superhet subject to additional image noise; one can thus set $T_{\text{image}}=0$. (ii) Even when monitoring (with the cooling LNA disconnected), the cavity mode is continuously cooled (albeit only slightly) by the receiver’s front-end LNA, connected to the cavity’s monitoring port via a lossy stub tuner and cable (substantially reducing the size of the cooling effect) –see main text. The ambient/reference mode temperature in the above formula, namely $T_{\text{mode}}^{0}$, should thus be set equal (again, see main text) to $T_{\text{mode}}^{\text{ambient}}\approx$ 255.2 K, somewhat less than 290 K. (iii) the ambient (reference) and post-cooled measurements are performed with the cavity under the same coupling conditions, namely with a single, critically-coupled monitoring port; the reflection amplitude off this port is always zero (during measurement of the noise power); thus, one can set $\Gamma_{c}=\Gamma_{c}^{0}=0$. These largely simplifying modifications lead to:
@@
$\Delta P=10\log_{10}\frac{{G_{\text{LNA}}[(T_{\text{min}}+T_{\text{mode}})+4T_{0}\frac{R_{\text{n}}}{Z_{0}}\frac{\lvert\Gamma_{\text{opt}}\rvert^{2}}{\lvert 1+\Gamma_{\text{opt}}\rvert^{2}}]+T_{\text{REC}}}}{{G_{\text{LNA}}[(T_{\text{min}}+T_{\text{mode}}^{\text{ambient}})+4T_{0}\frac{R_{\text{n}}}{Z_{\text{0}}}\frac{\lvert\Gamma_{\text{opt}}\rvert^{2}}{\lvert 1+\Gamma_{\text{opt}}\rvert^{2}}]+T_{\text{REC}}}},$ (17)
@@
where $T_{\text{mode}}$ is the pre-cooled mode temperature; $T_{\text{min}}$, $R_{\text{n}}$, $\Gamma_{\text{opt}}$ are the four noise parameters (being a complex number, $\Gamma_{\text{opt}}$ counts double) of the monitoring LNA, namely a Qorvo QPL9547EVB-01; the reference impedance $Z_{0}=50$ $\Omega$ throughout. The values assigned to these parameters, as shown in Table 1, have been interpolated to 1.45 GHz from measurements at straddling spot frequencies found in the manufacturer’s data sheet [49]. The noise temperature of the rest of the receiver, treated as a cascade of amplifiers, was calculated using the Friis formula and the amplifiers’ specified noise temperatures/figures [47]; see Table 1. The monitoring LNA’s gain, $G_{\text{LNA}}$, was measured using a VNA and found to be consistent with the QPL9547’s datasheet. The amplitude of the applied interrogating microwave tone was chosen to be sufficiently low that stimulated emission/absorption is very slow compared to the intrinsic spin dynamics of pentacene’s optically excited triplet state.
@@
Figure 14: Anatomy of the signal path at the front end of the homodyne receiver, displaying the forward and backward noise powers between the LNA and microwave cavity.
Table 1: Values of parameters -entering the noise-reduction equation 17
@@
Optimum source reflection coefficient $\Gamma_{\text{opt}}$ (0.073+0.125)i*
Noise resistance of first LNA $R_{\text{n}}$ 2.00*
@@
* interpolated from published datasheet [49]
$\dagger$ calculated using Friis formula[47]
@@
Given the parameters appearing in Table 1 whose values compose the noise calibration of our tr-EPR spectrometer, the observed reduction in noise power provided by cavity precooling, as predicted by Eq. 17, depends on the value $T_{\text{mode}}$ in the numerator versus that of $T_{\text{mode}}^{\text{ambient}}$ in the denominator. The latter, corresponding to when the cavity’s $\textrm{TE}_{01\delta}$ mode is critically coupled to its monitoring LNA via a lossy stub tuner, should equal $(T_{0}+T_{\text{mon}})/2=255.4$ K, where $T_{\text{mon}}=10^{-\frac{L_{\text{dB,2}}}{10}}T_{\text{cold}}+(1-10^{-\frac{L_{\text{dB,2}}}{10}})T_{0}\approx 220.9$ K according to Eq. 14. Upon fixing the value of $T_{\text{mode}}^{\text{ambient}}$, Eq. 17 can be used to infer $T_{\text{mode}}$ from the experimentally observed reduction in noise power $\Delta P$. Using the data shown in Fig. 6(b), a rough estimate of $\Delta P$ can be arrived at from the difference in the spectral noise power, averaged from 5 to 10 MHz (so avoiding the 1/f noise prevalent below 1 MHz), between “pre-cooled” and “ambient” conditions. one obtains: $\Delta P_{\text{exp-EPR}}=-3.7\pm 1$ dB. Since this estimate of $\Delta P$ has greater uncertainty than that derived directly from monitoring the pre-cooling in the time domain, namely $\Delta P_{\text{exp-Cool}}$, it is not further used.
@@
The mode temperature $T_{\text{mode}}$ can be separately determined by Eq.15; it depends on the (presumed identical) noise temperature of the two LNAs, $T_{\text{cold}}$, the coupling factor of the over-coupling loop $\kappa_{\text{over}}$, and the insertion losses of the RF switch $L_{\text{dB, 1}}$ and the monitoring LNA’s connection to the cavity line $L_{\text{dB,2}}$. One arrives at
@@
$T_{\text{mode, EPR}}=\frac{T_{\text{0}}+\kappa_{\text{over}}[(1-\frac{L_{\text{dB,1}}}{4.34})T_{\text{cold}}+\frac{L_{\text{dB,1}}}{4.34}T_{0}]+[10^{-\frac{L_{\text{dB,2}}}{10}}T_{\text{cold}}+(1-10^{-\frac{L_{\text{dB,2}}}{10}})T_{0}]}{2+\kappa_{\text{over}}}\approx 108.1\pm 5~{}\text{K}.$ (18)
@@
The values of $L_{\text{dB,1}}$ and $L_{\text{dB,2}}$ at 1.4495 GHz were measured using our VNA (HP-8753C) to be $0.19\pm 0.01$ dB and $6.05\pm 0.01$ dB, respectively. The coupling factor between the cavity and the cold load was characterized by the method(s) of Kajfez and Hwan [34] using the same VNA; we obtained $\kappa_{\text{over}}=3.8\pm 0.3$. The value of $T_{\text{cold}}$ is calculated to be $18.4\pm 0.5$ K based on the definition of noise parameters[46], utilizing the parameters listed in Table 1. $T_{0}$ represents room temperature, set at $290$ K. The uncertainty of $T_{\text{mode, EPR}}$ shown in Eq. 18 is calculated by applying Pythagoras’ theorem to error propagation. Based on this result, we can use Eq 17 to obtain $\Delta P_{\text{sim}}=-3.5\pm 0.2$ dB.

### v2

@@ II.1 Number of photons inside a microwave cavity
Figure S3: The active pre-cooling technique involves electromagnetically coupling a resonator, that is physically heat-sunk to a room-temperature environment (here referred to as the “hot bath”), to an active cold noise source or “load”, here functioning as a “cold bath”.
II.2 Mode temperature inside the cavity
@@
Upon achieving dynamic equilibrium, $\dot{q}=0$, the rate equation in Eq. S1 can thereupon be used to calculate the expected photon number. Defining the mode’s equivalent temperature as $T_{\text{mode}}=\frac{q}{\epsilon}=\frac{qhf_{\text{mode}}}{k_{\text{B}}}$, we obtain, for a single coupling port:
$T_{\text{mode}}=\frac{T_{0}+\kappa T_{\text{cold}}}{1+\kappa}.$ (S3)
@@
Note that this steady-state solution of the photon-number dynamics is consistent with the application of Siegman’s “useful noise theorem” (see chapter 8 of Ref. Siegman (1964)) as presented in the main text. In words: the mode’s expected temperature is the weighted average temperature of the two baths with which the mode interacts, the weighting of each bath being proportional to its interaction strength with the mode, i.e. its “coupling”. Here, the coupling of the mode to the cavity’s own intrinsic heat bath is (by definition/convention) unity. Fig. S4 displays the mode temperature predicted by Eq. S3. When the coupling factor of the over-coupling loop $\kappa$ is large, the mode temperature $T_{\text{mode}}$ will approach that of the cold load $T_{\text{cold}}$.
@@
Figure S4: Modeled equivalent temperature of an electromagnetic mode inside the cavity as a function of the coupling factor $\kappa$ (horizontal coordinate) and the noise temperature of the cold load $T_{\text{cold}}$ coupled to this mode (vertical coordinate).
II.3 Cavity cooling via a lossy transmission line
@@
The input of a low-noise amplifier (LNA) can serve as an active cold load for reducing the thermal noise inside a cavity that is coupled to it. Here we analyse the additional thermal noise generated by a lossy transmission line, such as a length of coaxial cable, or any other lossy component inserted between the cold load and the cavity, as shown generically in Fig. S5. Again, see chapter 8 of Ref. Siegman (1964) for the conceptual foundations and mathematics of the “wave approach”Dorđević et al. (2015) used here.
@@
Figure S5: Calculating how the cooling LNA’s noise temperature is degraded by transport along a length of (warm) lossy transmission line.
@@
$P_{l}=LP_{\text{cold}}+(1-L)P_{\text{line}},$ (S4)
@@
$P_{l}=(1-\frac{L_{\text{dB}}}{4.34})P_{\text{cold}}+\frac{L_{\text{dB}}}{4.34}P_{\text{line}}.$ (S5)
Note that this (slightly more complicated) approximation is more accurate than equation (8-2-5) in Siegman’s book Siegman (1964). Based on the relation between noise power and noise temperature, namely $P=k_{\text{B}}TB$, the effective noise temperature of the cold load, upon including the additional noise from a lossy transmission line connected to it, can thereupon be approximated as:
$T_{l}=(1-\frac{L_{\text{dB}}}{4.34})T_{\text{cold}}+\frac{L_{\text{dB}}}{4.34}T_{0},$ (S6)
where $T_{\text{cold}}$ the temperature of the cold load itself and $T_{0}=T_{\text{line}}$ is the (here assumed to be room) temperature of the transmission line. Combining Eq. S3 and Eq. S6, the mode’s noise temperature inside the microwave cavity is now given by:
$T^{\prime}_{\text{mode}}=\frac{T_{0}+\kappa[(1-\frac{L_{\text{dB}}}{4.34})T_{\text{cold}}+\frac{L_{\text{dB}}}{4.34}T_{0}]}{1+\kappa}=T_{\text{mode}}^{\text{orig.}}+\Delta T_{\text{link}},$ (S7)
where $T_{\text{mode}}^{\text{orig.}}$ is the original mode temperature as given by Eq. S3 above and $\Delta T_{\text{link}}=\frac{L_{\text{dB}}}{4.34}\kappa(T_{0}-T_{\text{cold}})/(1+\kappa)$ is the added noise temperature caused by the lossy transmission line.
II.4 Multiple cooling ports to a microwave cavity
We now consider a microwave cavity with two ports, one for cooling, the other for read-out, connected to separate LNAs functioning as cold loads. With the setup depicted in Fig. S6 and applying Eq. S2, the rate equation governing the microwave cavity’s photon number is:
@@ II.4 Multiple cooling ports to a microwave cavity
where $\kappa_{over}$ and $\kappa_{crit.}$ ($\kappa_{crit.}\equiv 1)$ are the coupling factors of the over-coupled and critically-coupled loops, respectively. As in the previous analysis for a single port, the depth of cavity cooling upon reaching dynamic equilibrium (i.e. “steady state” conditions) is deduced by solving $\dot{q}=0$. The mode temperature $T_{\text{mode}}$ achieved, as appears in Fig. S6, is given by:
$T_{\text{mode}}=\frac{T_{0}+(\kappa_{\text{over}}+\kappa_{\text{crit.}})T_{\text{cold}}}{1+\kappa_{\text{over}}+\kappa_{\text{crit.}}}=\frac{T_{0}+(1+\kappa_{\text{over}})T_{\text{cold}}}{2+\kappa_{\text{over}}}.$ (S9)
@@
$T_{\text{mode}}=\frac{T_{0}+\sum_{i=1}^{n}\kappa_{i}T_{i}}{1+\sum_{i=1}^{n}\kappa_{i}},$ (S10)
@@
Figure S6: Cavity cooling with two separate cold baths.
II.5 Multiple cooling ports to a microwave cavity with lossy transmission lines
In this section, we consider a microwave cavity with multiple cooling ports and lossy transmission lines, integrating concepts from sections 2.3 and 2.4. Combining equations Eq. S6 and Eq. S10, yields
$T_{\text{mode}}=\frac{T_{0}+\kappa_{1}[(1-\frac{L_{\text{dB,1}}}{4.34})T_{\text{cold}}+\frac{L_{\text{dB,1}}}{4.34}T_{0}]+\kappa_{2}[(1-\frac{L_{\text{dB,2}}}{4.34})T_{\text{cold}}+\frac{L_{\text{dB,2}}}{4.34}T_{0}]}{1+\kappa_{1}+\kappa_{2}},$ (S11)
where $\kappa_{1}$ and $\kappa_{2}$ represent the coupling coefficients of the first and second cold baths, respectively. We here assume the insertion losses of the transmission lines associated with the cold baths ($L_{\text{dB,1}}$, $L_{\text{dB,2}}$) are small. If however the insertion loss of a connecting cable/device is not small, one needs to revert back to Eq. S4, equivalent to equation in (8-2-4) in Siegman’s book Siegman (1964). Based on the relation between noise power and noise temperature, $P=k_{\text{B}}TB$, the equivalent noise temperature of a cold load at temperature $T_{\text{cold}}$, observed through a length of transmission line at temperature $T_{0}$ and of significant linear insertion loss $L=10^{-L_{\text{dB}/10}}$, can be quantified as:
$\begin{split}T_{l}&=LT_{\text{cold}}+(1-L)T_{0}\\ &=10^{-\frac{L_{\text{dB}}}{10}}T_{\text{cold}}+\left(1-10^{-\frac{L_{\text{dB}}}{10}}\right)T_{0}.\end{split}$ (S12)
@@ II.5 Multiple cooling ports to a microwave cavity with lossy transmission lines
$T_{\text{mode}}=\frac{T_{0}+\kappa_{1}[(1-\frac{L_{\text{dB,1}}}{4.34})T_{\text{cold}}+\frac{L_{\text{dB,1}}}{4.34}T_{0}]+\kappa_{2}[10^{-\frac{L_{\text{dB,2}}}{10}}T_{\text{cold}}+(1-10^{-\frac{L_{\text{dB,2}}}{10}})T_{0}]}{1+\kappa_{1}+\kappa_{2}}.$ (S13)
These approximations are analyzed in Fig. S7 below.
@@
Figure S7: How, according to Eq. S12, the forward noise power (in units of absolute temperature $T$) at the far end of a lossy cable depends on the cable’s insertion loss $L_{\text{dB}}$. Here the cable is uniformly maintained at a temperature of $T_{\text{0}}=290$ K throughout its length and the near end of the cable is terminated and cooled by the ACL whose noise temperature is $T_{\text{cold}}=29.1$ K (dashed cornflower-blue baseline)–corresponding to the input noise temperature of the Qorvo QPL9547EVB-01 LNA used experimentally. The zoomed-in area details how, when the insertion loss is less than 0.5 dB, equation Eq. S6 (red dashed line -“Linear Approx.”), i.e. the linear approximation provided by equation (8-2-5) in Siegman’s book Siegman (1964), compares against the exact formula Eq. S12 (blue solid line).
III Noise Analysis
@@
Similar to the previous work by Wu et al. (Wu et al., 2021), we also adopt the “wave approach” (see Chapter 8 of Siegman (Siegman, 1964)) to analyze the flow of noise power from the cavity into our homodyne receiver, as shown in Fig. S8. Equation S8 in the supplemental material of Wu et al.’s paper provides a formula for the factor, $\Delta P$ (in units of dB), by which the measured noise power is reduced by mode cooling:
@@
$\Delta P^{\prime}=10\log_{10}\frac{{G_{\text{LNA}}[(T_{\text{min}}+T_{\text{mode}})(1-|\Gamma_{c}|^{2})+4T_{0}\frac{R_{\text{n}}}{Z_{0}}\frac{\lvert\Gamma_{\text{c}}-\Gamma_{\text{opt}}\rvert^{2}}{\lvert 1+\Gamma_{\text{opt}}\rvert^{2}}+T_{\text{image}}]+T_{\text{REC}}}}{{G_{\text{LNA}}[(T_{\text{min}}+T_{\text{mode}}^{0})(1-|\Gamma_{c}^{0}|^{2})+4T_{0}\frac{R_{\text{n}}}{Z_{\text{0}}}\frac{\lvert\Gamma_{c}^{0}-\Gamma_{\text{opt}}\rvert^{2}}{\lvert 1+\Gamma_{\text{opt}}\rvert^{2}}+T_{\text{image}}]+T_{\text{REC}}}},$ (S14)
@@
However, this formula needs to be modified to suit our experimental set-up, and for the following reasons: (i) Our current receiver is a homodyne, not a single-conversion superhet subject to additional image noise; one can thus set $T_{\text{image}}=0$. (ii) Even when monitoring (with the cooling LNA disconnected), the cavity mode is continuously cooled (albeit only slightly) by the receiver’s front-end LNA, connected to the cavity’s monitoring port via a lossy stub tuner and cable (substantially reducing the size of the the LNA’s cooling effect) –see main text. The ambient/reference mode temperature in the above formula, namely $T_{\text{mode}}^{0}$, should thus be set equal (again, see main text) to $T_{\text{mode}}^{\text{ambient}}\approx$ 257.6 K, somewhat less than 290 K. Finally, (iii) the ambient (reference) and post-cooled measurements are performed with the cavity under the same coupling conditions, namely with a single, critically-coupled monitoring port; the reflection amplitude off this port is always zero (during measurement of the noise power); thus, one can set $\Gamma_{c}=\Gamma_{c}^{0}=0$. These largely simplifying modifications lead to:
@@
$\Delta P=10\log_{10}\frac{{G_{\text{LNA}}[(T_{\text{min}}+T_{\text{mode}})+4T_{0}\frac{R_{\text{n}}}{Z_{0}}\frac{\lvert\Gamma_{\text{opt}}\rvert^{2}}{\lvert 1+\Gamma_{\text{opt}}\rvert^{2}}]+T_{\text{REC}}}}{{G_{\text{LNA}}[(T_{\text{min}}+T_{\text{mode}}^{\text{ambient}})+4T_{0}\frac{R_{\text{n}}}{Z_{\text{0}}}\frac{\lvert\Gamma_{\text{opt}}\rvert^{2}}{\lvert 1+\Gamma_{\text{opt}}\rvert^{2}}]+T_{\text{REC}}}},$ (S15)
@@
where $T_{\text{mode}}$ is the pre-cooled mode temperature; $T_{\text{min}}$, $R_{\text{n}}$, $\Gamma_{\text{opt}}$ are the four noise parameters (being a complex number, $\Gamma_{\text{opt}}$ counts double) of the monitoring LNA, namely a Qorvo QPL9547EVB-01; the reference impedance $Z_{0}=50$ $\Omega$ throughout. The values assigned to these parameters, as shown in Table S1, have been interpolated to 1.45 GHz from measurements at straddling spot frequencies found in the manufacturer’s data sheet Qorvo (2023). The noise temperature of the rest of the receiver, treated as a cascade of amplifiers, was calculated using Friis’ formula and amplifiers’ specified noise temperatures and/or equivalent noise figures; see Table S1. The monitoring LNA’s gain, $G_{\text{LNA}}$, was measured using a VNA and found to be consistent with the QPL9547’s datasheet. The amplitude of the applied interrogating microwave tone was chosen to be sufficiently low that stimulated emission/absorption is very slow compared to the intrinsic spin dynamics of pentacene’s optically excited triplet state.
@@
Figure S8: Anatomy of the signal path at the front end of the homodyne receiver, displaying the forward and backward noise powers between the LNA and microwave cavity.
Table S1: Values of parameters -entering the noise-reduction equation S15
@@
Optimum source reflection coefficient $\Gamma_{\text{opt}}$ (0.073+0.125i)*
Noise resistance of first LNA $R_{\text{n}}$ 2.00 $\Omega$*
Reflection looking into input of first LNA $S_{11}$(0.107+0.234i)*
@@
* interpolated from published datasheet Qorvo (2023)
$\dagger$ calculated using Friis’ formula Friis (1944)
@@
Given the parameters appearing in Table S1, whose values compose the noise calibration of our tr-EPR spectrometer, the observed reduction in noise power provided by precooling the cavity, as predicted by Eq. S15, depends on the value $T_{\text{mode}}$ in the numerator versus that of $T_{\text{mode}}^{\text{ambient}}$ in the denominator. The latter, corresponding to when the cavity’s $\textrm{TE}_{01\delta}$ mode is critically coupled to its monitoring LNA via a lossy stub tuner, should equal $(T_{0}+T_{\text{mon.}})/2=257.6$ K, where $T_{\text{mon.}}=10^{-\frac{L_{\text{dB,2}}}{10}}T_{\text{cold}}+(1-10^{-\frac{L_{\text{dB,2}}}{10}})T_{0}\approx 225.2$ K according to Eq. S12. Upon fixing the value of $T_{\text{mode}}^{\text{ambient}}$, Eq. S15 can be used to infer $T_{\text{mode}}$ from the experimentally observed reduction in noise power $\Delta P$. Using the data shown in Fig. 8(b), a rough estimate of $\Delta P$ can be arrived at from the difference in the spectral noise power, averaged from 5 to 10 MHz (so avoiding the 1/f noise prevalent below 1 MHz), between “pre-cooled” and “ambient” conditions; one obtains: $\Delta P_{\text{exp-EPR}}=-3.7\pm 1$ dB. Since this estimate of $\Delta P$ has greater uncertainty than that derived directly from monitoring the pre-cooling in the time domain, namely $\Delta P_{\text{exp-Cool}}$, it is not further used.
@@
The mode temperature $T_{\text{mode}}$ can be separately predicted using our model, namely Eq. S13 above; it depends on the (presumed identical) noise temperature of the two LNAs, $T_{\text{cold}}$, the coupling factor of the over-coupling loop $\kappa_{\text{over}}$, and the insertion losses of the RF switch $L_{\text{dB, 1}}$ and the monitoring LNA’s connection to the cavity line $L_{\text{dB,2}}$. One arrives at
@@
$T_{\text{mode}}=\frac{T_{\text{0}}+\kappa^{\prime}_{\text{over}}[(1-\frac{L_{\text{dB,1}}}{4.34})T_{\text{cold}}+\frac{L_{\text{dB,1}}}{4.34}T_{0}]+[10^{-\frac{L_{\text{dB,2}}}{10}}T_{\text{cold}}+(1-10^{-\frac{L_{\text{dB,2}}}{10}})T_{0}]}{2+\kappa^{\prime}_{\text{over}}}\approx 131.5\pm 5~{}\text{K}.$ (S16)
@@
The values of $L_{\text{dB,1}}$ and $L_{\text{dB,2}}$ at 1.4495 GHz were measured using our VNA (HP-8753C) to be $0.19\pm 0.01$ dB and $6.05\pm 0.01$ dB, respectively. The coupling factor between the cavity and the cold load was characterized by the method(s) of Kajfez and Hwan Kajfez and Hwan (1984) using the same VNA; we obtained $\kappa_{\text{over}}=3.8\pm 0.3$. The value of $T_{\text{cold}}$ is calculated to be $29.1\pm 0.5$ K based on the definition of noise parameters Escotte et al. (1993); Escotte (2015) utilizing the parameters listed in Table S1. $T_{0}$ represents room temperature, set at $290$ K. The uncertainty of $T_{\text{mode}}$ shown in Eq. S16 is calculated by applying Pythagoras’ theorem to error propagation. Based on this result, we can use Eq. S15 to obtain $\Delta P_{\text{model}}=-3.2\pm 0.2$ dB.